OpenStudy (anonymous):

Write an equation in y intercept form of a line that passes through the points (-2,5) and (4,7)

OpenStudy (anonymous):

You have to do this first to find the slope\[\frac{ y^2-y^1 }{ x^2-x^1} \]

OpenStudy (anonymous):

The slope is 2/9

OpenStudy (anonymous):

How did u get a 9?

OpenStudy (anonymous):

Oh wait nevermind, its 2/6 :)

OpenStudy (anonymous):

I think I got the answer but I'm not sure. Is it y=2/6x+17/3 ?

OpenStudy (anonymous):

ok thats better! :) Next you have to find the equation y=mx+b First plug in the slope for m Then plug in on of the points for x and y Which will make you be able to find b

OpenStudy (anonymous):

I did not get the answer you have

OpenStudy (anonymous):

I checked it again. Thats what I really got

OpenStudy (anonymous):

I'm confident thats my answer

OpenStudy (anonymous):

Try this\[5=-\frac{ 2 }{ 6}(-2)+b\]

OpenStudy (anonymous):

The slope isnt -2/6 its 2/6

OpenStudy (anonymous):

Opps the negative is suppose to be there

OpenStudy (anonymous):

The negative IS suppose to be there? in the 2/6 ?

OpenStudy (anonymous):

This is what you should use\[5=\frac{ 2 }{ 6 }(-2)+b\]

OpenStudy (anonymous):

Yeah, that's what I did. After, I got 5=2/6(-2)+b and then I added 2/3 to both side and got b=17/3

OpenStudy (anonymous):

Was my answer right?

OpenStudy (anonymous):

b should be\[5\frac{ 2 }{ 3 }\] Oh i see what you did, You kept it as a mixed fraction, but yes thats right

OpenStudy (anonymous):

ah okay thanks:)

OpenStudy (anonymous):

Your Welcome!! :)